Two nuclides represented as (70 32X) and (37 17Y) are added to a proton to form anew nuclide. Determine the ratio of a proton-neutron of the new nuclide
1
Expert's answer
2021-08-24T06:52:28-0400

"_{32}^{70}X+_{17}^{37}Y+_1^0p=_{50}^{107}Sn,"

"Z(p)=50,"

"N(n)=107-50=57."